OpenStudy (itrymath):

2/3 (x - 2) = 4x i've never done 2/3 only 1/3

OpenStudy (itrymath):

@iGreen

OpenStudy (danieldjpon3):

what do you need help with???

OpenStudy (igreen):

Just distribute it.

OpenStudy (itrymath):

two thirds (x - 2) = 4x

OpenStudy (3mar):

\[\frac{ 2 }{ 3 }(x-2)=4x\]Is this your equation?

OpenStudy (igreen):

2/3(x - 2) 2/3 * x = ? 2/3 * -2 = ?

OpenStudy (itrymath):

what 3mar said

OpenStudy (itrymath):

2/3 (x - 2) = 4x

OpenStudy (3mar):

Ok. I am correct. Let's solve.

OpenStudy (itrymath):

how

OpenStudy (itrymath):

do i times the 3 * x?

OpenStudy (3mar):

Let's distribute and simplify. \[\frac{ 3 }{ 2 }(x-2)=\] multiply the fraction \[\frac{ 3 }{ 2 }\]by \[(x-2)\] we get:

OpenStudy (igreen):

2/3*

OpenStudy (igreen):

It's 2/3 * x And you just put the variable next to 2/3.

OpenStudy (igreen):

What about 2/3 * -2, can you multiply that?

OpenStudy (itrymath):

um hold on let me use calculator

OpenStudy (itrymath):

-1.333333333333333

OpenStudy (igreen):

Just multiply -2 to the numerator, 2 and keep the denominator the same. What do you get?

OpenStudy (3mar):

\[\frac{ 2 }{ 3 }x-\frac{ 2 }{ 3 }(-2)=\frac{ 2 }{ 3 }x-\frac{ 4 }{ 3 }\]

OpenStudy (itrymath):

what do 3mar do

OpenStudy (itrymath):

and i got hold on

OpenStudy (itrymath):

-4/6

OpenStudy (itrymath):

Thats what i got do i simplify

OpenStudy (igreen):

-4 is correct..but you don't multiply it to the denominator.

OpenStudy (itrymath):

omg so whats the answer -4/3

OpenStudy (igreen):

So we have: 2/3x - 4/3 = 4x Now subtract 2/3x to both sides, can you do that?

OpenStudy (3mar):

\[\frac{ 2 }{ 3 }x-\frac{ 4 }{ 3 }=4x\] Coolect all x in one side \[4x-\frac{ 2 }{ 3 }x=-\frac{ 4 }{ 3 }\] subtract: \[4-\frac{ 2 }{ 3 }=\frac{ 10 }{ 3 }\] then \[\frac{ 10 }{ 3 }x=-\frac{ 4 }{ 3 }\] so \[x=\frac{ -4 }{10 }=\frac{ 2 }{ 5 }=0.4\]

OpenStudy (3mar):

Hope that helps

OpenStudy (itrymath):

1/6?

OpenStudy (itrymath):

thanks 3mar but i like to be guided through the steps

OpenStudy (itrymath):

really do appreciate it

OpenStudy (itrymath):

i got 1/6 iGreen

OpenStudy (itrymath):

heello?

OpenStudy (itrymath):

@kittiwitti1 helllp

OpenStudy (itrymath):

@xapproachesinfinity help anyone

OpenStudy (3mar):

I am here. Where you stuck?

OpenStudy (itrymath):

how did you get 10/3?

OpenStudy (itrymath):

4*2 is 8 not 10

OpenStudy (3mar):

When you subtract the coefficients of the variable x, what are you doing? specially when there are fractions?

OpenStudy (3mar):

\[4-\frac{ 2 }{ 3 }=???\]

OpenStudy (itrymath):

what are coefficieants i forgot ?

OpenStudy (itrymath):

so 4*3 = 12-2 and we get 10?

OpenStudy (igreen):

@ItryMath No..how did you get 1/6?

OpenStudy (igreen):

It's 4 - 2/3

OpenStudy (igreen):

First try to get 4 to have a denominator of 3.

OpenStudy (3mar):